OpenStudy (anonymous):

\frac{ 2 }{ 3 }xge -6 Can you solve the equation?

OpenStudy (agent0smith):

assuming it's \[\frac{ 2 }{ 3 }x \ge -6 \] Try multiplying both sides by 3, then divide both sides by 2. Or multiply it all by 3/2: \[\frac{ 3 }{2 }\frac{ 2 }{ 3 }x \ge \frac{ 3 }{2} \times (-6)\]